A customer has asked a caterer for 60lb of nuts, 60% of which are to be cashews. The caterer have available mixtures of 70% cashews and 45% cashews. How many pounds of each mixture should be used?
Answer by Paul(988)

70x+45(60-x)=60(60)
70x+2700-45x=3600
25x=900
x=36

60-36=24

Hence, for 70% there are 36Lb of nuts, and for 45% there are 24lb of nuts.
Paul.